OpenStudy (anonymous):

Each dimension of a rectangular prism is increased by 10%. What percent of the original volume is the volume of the new prism?

OpenStudy (anonymous):

let the three dimension be x,y,z where original volume = x*y*z when all dimensions are increased by 10%, the new dimension of the prism will be 1.1x,1.1y,1.1z and thus volume of new prism = 1.1x * 1.1y * 1.1z = 1.331 * x *y* z So percent of the original volume is the volume of the new prism = xyz / ( 1.331 xyz) * 100% = 1/1.331 * 100% =75.13%

OpenStudy (anonymous):

These are my only possible answers.. 00.1 percent 133.1 percent 121 percent 1 percent

OpenStudy (anonymous):

oops , i placed the denominator and numerator in the wrong place sorry :) So percent of the original volume is the volume of the new prism = ( 1.331 xyz) / xyz * 100% = 1.331 * 100% =133.1%